OpenStudy (anonymous):

A recipe for molding dough calls for flour and salt in a ratio of 2 to 3. If you use 6 cups of salt, how much flour should you use?

OpenStudy (anonymous):

2/3 = x/6 x= 2*6/3 = 4 Cups of flour